# Track-On2: Enhancing Online Point Tracking with Memory

### v1

@@ Abstract
Long-term point tracking, Track any point, Online tracking, Synthetic data, Video understanding
@@ 1 Introduction
Figure 2: Offline vs. Online Point Tracking. We propose an online model, tracking points frame-by-frame (right), unlike the dominant offline paradigm where models require access to all frames within a sliding window or the entire video (left). To capture temporal information, we introduce a memory, storing contextual information from previous states of the point.
@@
Figure 3: Max GPU memory usage with increasing video length. We plot the maximum GPU memory usage of state-of-the-art models when tracking 256 points in videos of varying lengths. Models are grouped into three categories based on their input processing strategy: Video-level (CoTracker3 [11], BootsTAPIR [14]), Windowed (CoTracker3 with sliding windows), and Frame-level (TAPNext-B [12], Track-On2). It is observed that video-level models scale poorly and quickly run out of memory on longer sequences, while others remain efficient across all lengths.
@@ 2.1 Point Tracking
Fine-tuning on Real-world Videos: Due to the scarcity of manually annotated point tracks, most existing models are pretrained on the synthetic TAP-Vid-Kubric dataset [15]. Several recent efforts aim to improve real-world performance via domain adaptation: BootsTAPIR [14] uses student-teacher distillation on large-scale unlabeled video data, while CoTracker3 [11] further leverages pseudo-labeled real-world videos during training. Despite these advances, all such methods still rely heavily on synthetic pretraining, not only as initialization but also as a source of teacher models and supervision signals. Track-On2 focuses exclusively on synthetic supervision, systematically analyzing how to maximize its effectiveness. In particular, we identify training video length as the dominant factor shaping memory behavior and long-term generalization.
@@ 2.2 Causal Processing in Videos
Online or temporally causal models process frames sequentially, relying only on current and past inputs without future access. This contrasts with the dominant clip-based paradigm in point tracking. Causal processing is particularly useful for streaming [24, 25], embodied perception [26], and long videos [27, 28], where activation caching and scalability are important. As a result, causal modeling has been explored across tasks like pose estimation [29, 30], action detection [31, 32, 33, 34, 35, 36, 37, 28, 38], object tracking [39, 40], video captioning [25], and video object segmentation [41, 42].
Context propagation in causal models is typically handled via recurrent structures [32] or transformers with causal masks [36], but both face limitations in memory and temporal range. To improve efficiency and capacity, memory modules have been proposed. For example, LSTR [28] separates long- and short-term memory for action detection, and XMem [41] uses fine-grained memory in segmentation. Similar strategies extend to other domains like long-text processing [43] and video QA [44]. We also employ an attention-based memory mechanism.
@@ 3 Methodology
Figure 4: Overview. We introduce Track-On2, a simple transformer-based method for online, frame-by-frame point tracking. The process involves three steps: (i) Visual Encoder (bottom-left), which extracts multi-scale features from each frame using a DINOv3-based ViT-Adapter and fuses them through an FPN; (ii) Query Decoder, which decodes interest point queries by attending to the frame features and the memory propagated from the previous frame; (iii) Point Prediction (right), where correspondences are estimated in a coarse-to-fine manner, first through patch classification based on feature similarity, followed by offset prediction from the most likely patches, i.e. those most similar to the decoded queries. Before selecting the top patches, we perform re-ranking by enriching each query with local information from the top-$k$ candidate patches. After re-ranking, the refined queries are stored in memory for use in the next frame’s estimation. Note that $\Box$ refers to point query, while ⃝ represents prediction, either as point coordinate or visibility.
@@ 3.2.2 Query Decoder
Memory: The memory maintains a per-query history of visual embeddings for all $N$ queries. Each query has a slot that holds up to $L$ entries. As time advances, updates follow a FIFO policy: when full, the oldest embedding is evicted to accommodate the newest. This yields a bounded, adaptive temporal context whose size does not grow with video length.
@@
Figure 8: Re-ranking. In certain cases, a patch with high similarity, though not the most similar, is closer to the ground-truth patch (left). The top-33 patch centers, ranked by similarity, are marked with dots, while the ground-truth is represented by a diamond. To address such cases, we enrich the queries with local information from the top-ranked patches using the re-ranking module (right). Given the decoded queries $\mathbf{q}^{dec}_{t}$ and the similarity map $\mathbf{C}^{dec}_{t}$, we first identify the top-$k$ patches with the highest similarity scores. Their corresponding features are retrieved from the multi-scale features using a deformable cross-attention (Def. CA). The resulting top-$k$ features are then fused into the original queries via a cross-attention (CA) to obtain refined queries $\mathbf{q}_{t}$. Using $\mathbf{q}_{t}$, we compute the final similarity map for coarse location prediction.
@@ 3.2.3 Point Prediction
Re-ranking: We observed that the true target patch does not always achieve the highest score in $\mathbf{C}^{{dec}}_{t}$, but it typically appears within the top-k. For instance, Fig. 8(a) shows a case where the top-2 patch is closer to the true match than the top-1. To address this, we introduce a re-ranking module $\Phi_{\text{re-rank}}$:
@@ 4.1 Experimental Setup
Notably, in Track-On, the training video length and memory size were limited to 24 and 12, respectively. In contrast, Track-On2 is trained with longer sequences and expanded memory capacity. As demonstrated in the following comparisons and ablations, the interaction between these two factors plays a critical role in improving temporal reasoning and long-range tracking performance.
4.2 Results
@@
4.2.1 TAP-Vid DAVIS
Track-On2 sets the best results across all metrics. While the DINOv3 variant performs strongest, the closest competitor is still our own model with a DINOv2 backbone, differing only by small margins. Among online methods trained only on synthetic data, it improves over its predecessor Track-On by +2.0 AJ, +1.9 $\delta^{x}_{avg}$, and +1.2 OA. This boost, specifically on this dataset, is primarily attributed to training with longer video clips, which equips the model with more durable temporal representations (see Sec. 4.5.3). It further surpasses strong baselines fine-tuned on real videos: compared to BootsTAPNext-B [12], Track-On2 is +1.8 AJ (67.0 vs. 65.2), +1.4 $\delta^{x}_{avg}$(79.9 vs. 78.5), and +0.8 OA (92.0 vs. 91.2). These results are particularly impressive because our model is an online approach, processing the video frame by frame, and trained only on synthetic data, yet it exceeds the performance of offline models, fine-tuned on real-world data, that process the entire video at once, such as the video variant of CoTracker3 [11].
@@ 4.2.2 TAP-Vid Kinetics
Kinetics comprises long, diverse Internet videos. Within the synthetic-only group, Track-On2 outperforms TAPTRv3 [51] in AJ and OA (55.3 vs. 54.5 and 89.6 vs. 88.2) and exceeds TAPNext-B in $\delta^{x}_{avg}$(69.3 vs. 67.9). The difference between our default DINOv3 model and its DINOv2 counterpart remains minor, reinforcing that the gains primarily stem from architectural changes and training setup rather than the backbone alone. While real-world fine-tuning improves results across all models on this dataset, with $\delta^{x}_{avg}$gains of +2.8 for TAPNext-B, +1.9 for CoTracker3 (Window), and +1.3 for CoTracker3 (Video); Track-On2 remains highly competitive without any real-video supervision. It achieves the highest OA and remains close in AJ and $\delta^{x}_{avg}$to fine-tuned counterparts, despite relying solely on synthetic data. We also note the parameter efficiency: BootsTAPNext-B uses 194M parameters, whereas our full model has 52.3M parameters with only 23.6M learnable (thanks to frozen visual encoder), i.e. roughly one quarter of BootsTAPNext-B’s learnable capacity.
4.2.3 RoboTAP
@@
RoboTAP consists of robotic sequences with an average length of 274 frames, and our model consistently surpasses existing online and offline models across all metrics. Among synthetic-only methods, it surpasses TAPTRv3 by +3.5 AJ, +3.3 $\delta^{x}_{avg}$, and +3.3 OA. This demonstrates that our memory modules, which enable online tracking, are capable of effectively capturing the dynamics of long video sequences despite lacking bidirectional information flow across all frames. On this dataset, we also observe a clear benefit from using DINOv3 over DINOv2. While the DINOv2 variant already outperforms previous models in AJ and OA, the DINOv3 backbone further boosts performance. As with Kinetics, real-world fine-tuning brings substantial gains in RoboTAP: AJ increases by +5.6 and +4.8 for CoTracker3 (Window) and CoTracker3 (Video), respectively. Despite not using any real-video fine-tuning, Track-On2 still outperforms all real fine-tuned baselines by a clear margin. We emphasize that RoboTAP reflects a robotics setting where online operation is crucial; Track-On2’s strong results here indicate its suitability for real-time embodied perception and related tasks.
4.2.4 Dynamic Replica
@@
This dataset comprises 300-frame videos and is suitable to test long-term tracking. Although CoTracker3 reports the scores of other models on this dataset, their official implementation samples 256 tracks randomly from the available annotations. For fair comparison, we re-evaluated the available checkpoints using uniformly sampled tracks instead of random sampling. On this benchmark, both variants of Track-On2 surpass prior work. The DINOv2 version achieves the highest $\delta^{x}_{avg}$, outperforming CoTracker3 (Window) by +2.3. The default DINOv3 configuration also outperforms all previous models, but remains slightly behind our DINOv2 variant by 0.1 points. While BootsTAPIR and CoTracker3 remain competitive (68.4 and 72.3 $\delta^{x}_{avg}$), TAPNext and BootsTAPNext are considerably lower at 47.8 and 46.2. This drop is consistent with the limitation acknowledged in TAPNext, where predictions deteriorate once the sequence length exceeds a threshold, despite strong results on shorter clips. Overall, these results emphasize the importance of temporal robustness and generalization to long sequences; our memory enables stable tracking over hundreds of frames rather than being tied to a fixed clip length.
@@
4.2.5 PointOdyssey
PointOdyssey is specifically designed to evaluate tracking on very long videos, with an average length of 2,386 frames and a maximum of 4,325. This dataset is particularly important for testing temporal robustness and generalization, since existing models are trained on much shorter sequences: 64 frames for the CoTracker3 family; 48 frames for our model, and the TAPNext family; and 24 frames for Track-On, and BootsTAPIR. This distinction is especially critical for online models, as recovering from errors in fully causal tracking is harder, and mistakes accumulate more severely compared to shorter sequences. Despite this handicap, both variants of Track-On2 achieve the best performance in $\delta^{x}_{avg}$and survival rate (the fraction of tracks that remain valid without large errors), surpassing even the window-based variant of CoTracker3. Interestingly, the DINOv2 backbone proves more effective than DINOv3 on this dataset: it outperforms DINOv3 across all metrics and even achieves the lowest Median Trajectory Error (MTE) among all models.
@@ 4.2.6 Comparison to Track-On
Track-On2 consistently outperforms its predecessor Track-On across all benchmarks. It achieves gains of +2.0 AJ on TAP-Vid DAVIS, +2.3 AJ on Kinetics, and up to +4.6 AJ on RoboTAP, when the same DINOv2 backbone is considered. Most notably, the improvement on PointOdyssey is substantial, highlighting Track-On2’s robustness on extremely long sequences. While architectural refinements contribute to these gains, the primary factor is training on longer video clips, which allows the memory module to capture richer temporal context and generalize more effectively to long-term dependencies. These contributions are further analyzed in the ablation study (Sec. 4.5). In addition to improved performance, the architectural changes in Track-On2 also bring substantial efficiency gains (see Section 4.4), placing it in a strong position for real-world applications where streaming, frame-by-frame inference is essential.
@@ 4.3 Qualitative Results
Additionally, we present sparse tracking results on a RoboTAP video for three top-performing models, namely CoTracker3 (Window), BootsTAPNext, and our model, in Fig. 11. Query points are shown on the first frame. We observe that CoTracker3 fails to recover points on the green block once lost (third column), whereas both BootsTAPNext and our model can continue tracking them after occlusion. Later in the sequence, BootsTAPNext loses track of the cube (fifth column), leaving Track-On2 as the only model able to maintain these points until the end. As the video progresses, BootsTAPNext predictions diverge and are predicted occluded, consistent with the limitations reported for TAPNext. Interestingly, after a long static period, CoTracker3 keeps predicting cyan query points on the left robotic arm at their previous location as visible, even though the arm has moved (top row, last two columns). We speculate this is a side effect of its iterative updates, where the model becomes biased towards expected motion rather than appearance cues, assuming the point remains constant after extended inactivity.
Overall, these results underline the strengths of our memory design: Track-On2 preserves trajectories over long sequences, recovers from occlusions, and maintains stable predictions where competing methods fail. However, regions with repetitive or textureless appearance remain challenging, as points lack strong anchors for accurate localization.

### v2

@@ Abstract
Long-term point tracking, Track any point, Online tracking, Video understanding
@@ 1 Introduction
Figure 2: Offline vs. Online Point Tracking. We propose an online model, tracking points frame-by-frame (right), unlike the dominant offline paradigm where models require access to all frames within a sliding window or the entire video (left). To capture temporal information, we introduce a memory, storing contextual information from previous states of the point. The blue box denotes the initial query sampled from the first-frame point location, and the purple boxes represent the decoded queries at each timestep, conditioned on both the initial sample and the evolving memory.
@@
Figure 3: Max GPU memory usage with increasing video length. We plot the maximum GPU memory usage of state-of-the-art models when tracking 256 points in videos of varying lengths. Models are grouped into three categories based on their input processing strategy: Video, window, and frame. It is observed that video-level models scale poorly and quickly run out of memory on longer sequences, while others remain efficient across all lengths
@@ 2.1 Point Tracking
Fine-tuning on Real-world Videos: Due to the scarcity of manually annotated point tracks, most existing models are pretrained on the synthetic TAP-Vid-Kubric dataset [14]. Several recent efforts aim to improve real-world performance via domain adaptation: BootsTAPIR [21] uses student-teacher distillation on large-scale unlabeled video data, while CoTracker3 [11] further leverages pseudo-labeled real-world videos during training. Despite these advances, all such methods still rely heavily on synthetic pretraining, not only as initialization but also as a source of teacher models and supervision signals. Track-On2 focuses exclusively on synthetic supervision, systematically analyzing how to maximize its effectiveness.
@@ 2.2 Causal Processing in Videos
Online or temporally causal models process frames sequentially, relying only on current and past inputs without future access. This contrasts with the dominant clip-based paradigm in point tracking. Causal processing is particularly useful for streaming [24, 25], embodied perception [26], and long videos [27, 28], where activation caching and scalability are important. As a result, causal modeling has been explored across tasks like pose estimation [29, 30], action detection [31, 32, 33, 34, 35, 36, 37, 28, 38], object tracking [39, 40], video captioning [25], and video object segmentation [41, 42]. Context propagation in causal models is typically handled via recurrent structures [32] or transformers with causal masks [36], but both face limitations in memory and temporal range. To improve efficiency and capacity, memory modules have been proposed. For example, LSTR [28] separates long- and short-term memory for action detection, and XMem [41] uses fine-grained memory in segmentation. In visual tracking, ODTrack [43] reformulates tracking as online token propagation, enabling dense inter-frame association without image-pair matching. UM-ODTrack [44] extends this idea to a unified framework while maintaining temporally causal processing. Similar memory-augmented or token-based propagation strategies have also been explored in long-text modeling [45] and video question answering [46].
@@ 3 Methodology
Figure 4: Overview. We introduce Track-On2, a transformer-based method for online, frame-by-frame point tracking. The pipeline consists of three stages: (i) Visual Encoder (top-left), which extracts multi-scale features from each frame using a DINOv3-based ViT-Adapter and fuses them via an FPN; (ii) Query Decoding (bottom-left), where point queries attend to the current frame features and a persistent memory propagated from the previous frame; (iii) Point Prediction (right), which estimates correspondences in a coarse-to-fine manner. Decoded queries are refined by a re-ranking module that incorporates local information from candidate matches, and point locations are predicted by patch-level classification with sub-patch regression, alongside a lightweight visibility head. Refined queries are appended to memory for use in the next frame. $\Box$ denotes a point query, and ⃝ denotes a prediction.
@@ 3.2.2 Query Decoder
Memory: The memory maintains a per-query history of visual embeddings for all $N$ queries. Each query has a slot that holds up to $L$ entries. As time advances, updates follow a FIFO policy: when full, the oldest embedding is evicted to accommodate the newest. This yields a bounded, adaptive temporal context whose size does not grow with video length. This design provides a causal, bounded temporal context that favors recent observations, reducing error accumulation and improving robustness to abrupt changes such as long occlusions or fast motion, while avoiding the state overfitting and limited spatial flexibility commonly associated with recurrent memory [49].
@@
Figure 8: Re-ranking. In some cases, a patch with high similarity, though not ranked first, is spatially closer to the ground-truth patch (left). The top-33 patch centers ranked by similarity are shown as dots, while the ground truth is marked by a diamond. To address this, we refine decoded queries using a re-ranking module (right). Starting from $\mathbf{q}^{dec}_{t}$ and $\mathbf{C}^{dec}_{t}$, we select the top-$k$ patches, extract their multi-scale features via deformable cross-attention (Def. CA), and fuse them into the queries using cross-attention (CA) to obtain refined queries $\mathbf{q}_{t}$ for final similarity prediction.
@@ 3.2.3 Point Prediction
Re-ranking: We observed that the true target patch does not always achieve the highest score in $\mathbf{C}^{{dec}}_{t}$, but it typically appears within the top-k. For instance, Fig. 8(a) shows a case where the top-2 patch is closer to the true match than the top-1. We validate the premise of re-ranking via an oracle analysis on DAVIS in Fig. 9, measuring the distance from the ground truth to the closest of the top-$k$ predictions in the correlation map $\mathbf{C}$. As $k$ increases, this distance decreases consistently, indicating that accurate candidates are often present even when the top-1 match is incorrect. Rather than enforcing deterministic selection, using top-$k$ candidates increases the likelihood that the correct match is available for subsequent refinement. Motivated by this finding, we introduce a re-ranking module $\Phi_{\text{re-rank}}$:
@@
Figure 9: Re-ranking oracle. L2 distance to the ground truth versus the number of top-$k$ candidates from the correlation $\mathbf{C}$, measured by the closest candidate among the top-$k$.
@@ 4.1 Experimental Setup
4.2 Quantitative Results
@@
4.2.1 Comparison with Previous Work
TAP-Vid DAVIS: Track-On2 sets the best results across all metrics. While the DINOv3 variant performs strongest, the closest competitor is still our own model with a DINOv2 backbone, differing only by small margins. Among online methods trained only on synthetic data, it improves over its predecessor Track-On by +2.0 AJ, +1.9 $\delta^{x}_{avg}$, and +1.2 OA. This boost, specifically on this dataset, is primarily attributed to training with longer video clips, which equips the model with more durable temporal representations (see Sec. 4.5.4). It further surpasses strong baselines fine-tuned on real videos: compared to BootsTAPNext-B [12], Track-On2 is +1.8 AJ (67.0 vs. 65.2), +1.4 $\delta^{x}_{avg}$(79.9 vs. 78.5), and +0.8 OA (92.0 vs. 91.2). These results are particularly impressive because our model is an online approach, processing the video frame by frame, and trained only on synthetic data, yet it exceeds the performance of offline models, fine-tuned on real-world data, that process the entire video at once, such as the video variant of CoTracker3 [11].
@@ 4.2.1 Comparison with Previous Work
TAP-Vid Kinetics: Kinetics comprises long, diverse Internet videos. Within the synthetic-only group, Track-On2 outperforms TAPTRv3 [54] in AJ and OA (55.3 vs. 54.5 and 89.6 vs. 88.2) and exceeds TAPNext-B in $\delta^{x}_{avg}$(69.3 vs. 67.9). The difference between our default DINOv3 model and its DINOv2 counterpart remains minor, reinforcing that the gains primarily stem from architectural changes and training setup rather than the backbone alone. While real-world fine-tuning improves results across all models on this dataset, with $\delta^{x}_{avg}$gains of +2.8 for TAPNext-B, +1.9 for CoTracker3 (Window), and +1.3 for CoTracker3 (Video); Track-On2 remains highly competitive without any real-video supervision. It achieves the highest OA and remains close in AJ and $\delta^{x}_{avg}$to fine-tuned counterparts, despite relying solely on synthetic data. We also note the parameter efficiency: BootsTAPNext-B uses 194M parameters, whereas our full model has 52.3M parameters with only 23.6M learnable (thanks to frozen visual encoder), i.e. roughly one quarter of BootsTAPNext-B’s learnable capacity.
@@
RoboTAP: RoboTAP consists of robotic sequences with an average length of 274 frames, and our model consistently surpasses existing online and offline models across all metrics. Among synthetic-only methods, it surpasses TAPTRv3 by +3.5 AJ, +3.3 $\delta^{x}_{avg}$, and +3.3 OA. This demonstrates that our memory modules, which enable online tracking, are capable of effectively capturing the dynamics of long video sequences despite lacking bidirectional information flow across all frames. On this dataset, we also observe a clear benefit from using DINOv3 over DINOv2. While the DINOv2 variant already outperforms previous models in AJ and OA, the DINOv3 backbone further boosts performance. As with Kinetics, real-world fine-tuning brings substantial gains in RoboTAP: AJ increases by +5.6 and +4.8 for CoTracker3 (Window) and CoTracker3 (Video), respectively. Despite not using any real-video fine-tuning, Track-On2 still outperforms all real fine-tuned baselines by a clear margin. We emphasize that RoboTAP reflects a robotics setting where online operation is crucial; Track-On2’s strong results here indicate its suitability for real-time embodied perception and related tasks.
@@
Dynamic Replica: This dataset comprises 300-frame videos and is suitable to test long-term tracking. Although CoTracker3 reports the scores of other models on this dataset, their official implementation samples 256 tracks randomly from the available annotations. For fair comparison, we re-evaluated the available checkpoints using uniformly sampled tracks instead of random sampling. On this benchmark, both variants of Track-On2 surpass prior work. The DINOv2 version achieves the highest $\delta^{x}_{avg}$, outperforming CoTracker3 (Window) by +2.3. The default DINOv3 configuration also outperforms all previous models, but remains slightly behind our DINOv2 variant by 0.1 points. While BootsTAPIR and CoTracker3 remain competitive (68.4 and 72.3 $\delta^{x}_{avg}$), TAPNext and BootsTAPNext are considerably lower at 47.8 and 46.2. This drop is consistent with the limitation acknowledged in TAPNext, where predictions deteriorate once the sequence length exceeds a threshold, despite strong results on shorter clips. Overall, these results emphasize the importance of temporal robustness and generalization to long sequences; our memory enables stable tracking over hundreds of frames rather than being tied to a fixed clip length.
@@
PointOdyssey: PointOdyssey is specifically designed to evaluate tracking on very long videos, with an average length of 2,386 frames and a maximum of 4,325. This dataset is particularly important for testing temporal robustness and generalization, since existing models are trained on much shorter sequences: 64 frames for the CoTracker3 family; 48 frames for our model, and the TAPNext family; and 24 frames for Track-On, and BootsTAPIR. This distinction is especially critical for online models, as recovering from errors in fully causal tracking is harder, and mistakes accumulate more severely compared to shorter sequences. Despite this handicap, both variants of Track-On2 achieve the best performance in $\delta^{x}_{avg}$and survival rate (the fraction of tracks that remain valid without large errors), surpassing even the window-based variant of CoTracker3. Interestingly, the DINOv2 backbone proves more effective than DINOv3 on this dataset: it outperforms DINOv3 across all metrics and even achieves the lowest Median Trajectory Error (MTE) among all models.
@@ 4.2.2 Comparison to Track-On
Track-On2 consistently outperforms its predecessor Track-On across all benchmarks. It achieves gains of +2.0 AJ on TAP-Vid DAVIS, +2.3 AJ on Kinetics, and up to +4.6 AJ on RoboTAP, when the same DINOv2 backbone is considered. Most notably, the improvement on PointOdyssey is substantial, highlighting Track-On2’s robustness on extremely long sequences. While architectural refinements contribute to these gains, the primary factor is training on longer video clips, which allows the memory module to capture richer temporal context and generalize more effectively to long-term dependencies. These contributions are further analyzed in the ablation study (Sec. 4.5.1).
4.2.3 Analysis on Long-term Tracking
To further analyze stability without the confounding factor of cumulative failure, we present a frame-wise accuracy analysis in Fig. 13. Here, we report the percentage of points with error <16<16px at each timestep, allowing us to visualize how accuracy degrades over time, on Dynamic Replica and Kinetics. Track-On2 demonstrates the strongest overall robustness. It performs competitively with CoTracker3 on Kinetics; and on Dynamic Replica, our model outperforms all baselines. Notably, while methods like BootsTAPNext suffer sharp drops as sequence length increases, Track-On2 maintains a gentle decay slope, confirming its resistance to long-term drift compared to existing methods, including the original Track-On.
Figure 13: Long-term survival comparison. Survival rate as a function of frame number on Kinetics (left) and Dynamic Replica (right). The survival rate measures the percentage of points that remain successfully tracked over time.
@@ 4.3 Qualitative Results
Additionally, we present sparse tracking results on a RoboTAP video for three top-performing models, namely CoTracker3 (Window), BootsTAPNext, and our model, in Fig. 12. Query points are shown on the first frame. We observe that CoTracker3 fails to recover points on the green block once lost (third column), whereas both BootsTAPNext and our model can continue tracking them after occlusion. Later in the sequence, BootsTAPNext loses track of the cube (fifth column), leaving Track-On2 as the only model able to maintain these points until the end. As the video progresses, BootsTAPNext predictions diverge and are predicted occluded, consistent with the limitations reported for TAPNext. Interestingly, after a long static period, CoTracker3 keeps predicting cyan query points on the left robotic arm at their previous location as visible, even though the arm has moved (top row, last two columns). We speculate this is a side effect of its iterative updates, where the model becomes biased towards expected motion rather than appearance cues, assuming the point remains constant after extended inactivity. Overall, these results underline the strengths of our design: Track-On2 preserves trajectories over long sequences, recovers from occlusions, and maintains stable predictions where competing methods fail. However, regions with repetitive or textureless appearance remain challenging, as points lack strong anchors for accurate localization.